It took Lisa 2 hrs to read the first 100 pages of a book , if she reads the other 44 in an hour how many pages does she read per minute
Answer by CubeyThePenguin(3113)

Total time: 2 + 1 = 3 hours
Total pages read: 100 + 44 = 144 pages

reading speed = 144/3 = 48 pg/hr = 0.8 pg/min